# Isoparametric foliations and bounded geometry

### v1

@@ 1. Introduction
    The reason why dimension $n=5$ in Theorem 1.1 is special is that our proof relies on the fact that for a closed $(n-1)$-manifold $L$ the forgetful morphism $\pi_{0}(\operatorname{Diff}(L))\to\pi_{0}(\operatorname{Homeo}(L))$ has finite kernel for $n\neq 5$ (see Lemma 5.6). Here $\pi_{0}(\operatorname{Diff}(L))$ denotes the group of smooth isotopy classes of diffeomorphisms and $\pi_{0}(\operatorname{Homeo}(L))$ the group of topological isotopy classes of homeomorphisms. The analogous statement for $n=5$ is known to be false (see Remark 5.7).
@@
    A leaf is singular if it has codimension $>1$ and regular if it has codimension $1$.
@@ Acknowledgments
A.L. would like to thank the Hausdorff Research Institute of Mathematics in Bonn where this work was finalised during a stay of A.L. as part of the Trimester Program “Metric Analyis”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC-2047/1–390685813.
@@ 2.1. Transnormal functions, manifold submetries, and double disc bundles
The second alternative point of view is that of double disc bundles. To introduce it, we say that Euclidean disc bundle $D\rightarrow L$ is the unit disc bundle of a vector bundle equipped with a fiberwise metric over a closed manifold $L$ which may be disconnected and have components of different dimension. A double disc bundle decomposition of $M$ consists of two Euclidean disc bundles $D^{\pm}\rightarrow L^{\pm}$, a diffeomorphism $\varphi\colon\partial D^{+}\rightarrow\partial D^{-}$ between the boundaries of their total spaces, and a diffeomorphism $\tau\colon D^{+}\cup_{\varphi}D^{-}\rightarrow M$ from the manifold obtained by gluing $D^{\pm}$ along the boundaries using $\varphi$. Given such decomposition, there is a canonical map $P_{D^{\pm}}\colon M\rightarrow[-1,1]$ which is given as the composition of $\tau^{-1}$ with the map $D^{+}\cup_{\varphi}D^{-}$ that sends $e\in D^{\pm}$ to ±$\pm(\lvert e\rvert-1)$, using the fiberwise norm of $D^{\pm}$. The fibers over ±$\pm 1$ are the 00-sections $L^{\pm}\subset M$ and the fibers over $t\in(-1,1)$ are the concentric sphere subbundles of $D^{\pm}$.
@@
We first show that (ii) and (iii) are equivalent. Given a manifold submetry $P\colon(M,g)\rightarrow[c,d]$ as in (ii), we can without changing the fibers rescale the metrics so that $c=-1$ and $d=1$ (up to an opportune translation of the codomain). Now consider the smooth submanifolds $L^{\pm}\coloneqq P^{-1}(\pm 1)$ and the normal exponential map $\exp^{\pm}\colon\nu_{g}(L^{\pm})\rightarrow M$ from the normal bundle $\nu_{g}(L^{\pm})\subset TM$ of $L^{\pm}$ with respect to $g$. Using that $P$ is a submetry and that geodesics in $[-1,1]$ with length $<2$ starting from ±$\pm 1$ minimize distance, one sees that the restrictions $\exp^{\pm}\colon D_{g}(L^{\pm})\rightarrow M$ to the unit disc bundles $D_{g}(L^{\pm})\subset\nu_{g}(L^{\pm})$ are diffeomorphisms onto $P^{-1}([-1,0])$ and $P^{-1}([0,1])$ respectively (c.f. [KL22, Section 3.4]). Under this identification, the map $P\colon M\rightarrow[-1,1]$ corresponds to $D_{g}(L^{\pm})\rightarrow[-1,1],e\mapsto\pm(\lvert e\rvert-1)$, so we obtain a double disc bundle composition $\smash{D_{g}(L^{+})\cup_{\varphi}D_{g}(L^{-})\cong M}$ with $\smash{P_{D_{g}(L^{\pm})}=P}$ and gluing diffeomorphism $\varphi=(\exp^{-})^{-1}\circ\exp^{+}$ involving the diffeomorphism $\smash{\exp^{\pm}\colon\partial D_{g}(L^{\pm})\rightarrow P^{-1}(0)}$ given by restricting the normal exponential map to the unit sphere bundle. This shows that (ii) implies (iii). To show the converse, one constructs a suitable metric $g$ starting from a double disc bundle decomposition (see [GWY19, Lemma 3.1, Remark 3.2]).
@@
To prove that (i) and (ii) are equivalent, we use that the fibers of a transnormal function $f\colon M\rightarrow\mathbb{R}$ as in (i) are pairwise equidistant smooth submanifolds of $M$ by [Wan87]. This implies that there is a canonical metric on $f(M)$ that makes $f\colon M\rightarrow f(M)$ into a manifold submetry. By the mean value theorem, $f(M)$ is homeomorphic to a compact interval and thus also isometric to a compact interval with the standard metric. Postcomposing $f$ with this isometry yields a manifold submetry as in (ii) with the same fibers as $f$. Vice versa, given a manifold submetry $P:M\to[-c,d]$, we may assume $[c,d]=[-1,1]$ by rescaling the metric. We choose a smooth and strictly monotonically inreasing function $h\colon[-1,1]\rightarrow\mathbb{R}$ with $h(-1)=-1$, $h(1)=1$, $h^{\prime}(t)>0$ on $(-1,1)$, and such that $h(t)=(t+1)^{2}-1$ near $-1$ and $h(t)=1-(t-1)^{2}$ near $1$. The function $f\coloneq(h\circ P)$ has the same fibers as $P$, so it suffices to check that it is transnormal. As observed above, we have $P(-)=\pm 1\mp d_{M}(-,L^{\pm})$. Moreover, we have $(P\circ\exp^{\pm})(v)=\pm 1\mp|v|$ for $|v|\leq 2$, so it follows that $f=h\circ P$ is a smooth function. To check that it satisfies the first equation in (1.1), note that the differential of $f$ vanishes exactly on $L^{\pm}$. Near $L^{\pm}$ the function $f$ satisfies $\|\nabla f(\exp^{\pm}(v))\|^{2}=\|v\|^{2}$, which is a linear function of $f$. Outside of $L^{\pm}$, we have $\|\nabla f\|^{2}=\|h^{\prime}(P(x))\|^{2}=\|h^{\prime}\circ h^{-1}\circ f\|^{2}$. From this one concludes that $f$ satisfies indeed the first equation in (1.1), so it is transnormal.∎
@@ 2.2. Connectedness of fibers and existence of singular fibers
    $\mathcal{F}$ arises as an isoparametric foliation on $M$with respect to some Riemannian metric on $g$ in the sense of the introduction.
@@ 2.3. Foliated diffeomorphism, concordance, and isotopy
It is convenient to replace $D^{+}\cup_{\varphi_{i}}D^{-}$ up to foliated diffeomorphism with $\smash{D^{+}\cup_{\mathrm{id}_{\partial D^{+}}}\big{(}\partial D^{+}\times[0,1]\big{)}\cup_{\varphi_{i}}D^{-}}$. Here the latter is equipped with the partition into submanifolds given as the fibers of the map to $[-1,2]$ that sends $e\in D^{+}$ to $\lvert e\rvert-1$, $e\in D^{-}$ to $2-\lvert e\rvert$, and agrees with the projection to $[0,1]$ on the middle part. If the $\varphi_{i}$ are smoothly isotopic, then $\mathrm{id}_{\partial D^{+}}$ and $\varphi_{0}^{-1}\circ\varphi_{1}$ are smoothly isotopic. A choice of isotopy $\phi\colon\partial D^{+}\times[0,1]\to\partial D^{+}\times[0,1]$ between them induces a diffeomorphism
$(\mathrm{id}_{D^{+}}\cup\phi\cup\mathrm{id}_{D^{-}})\colon\Big{(}D^{+}\cup_{\mathrm{id}_{\partial D^{+}}}\big{(}\partial D^{+}\times[0,1]\big{)}\cup_{\varphi_{0}}D^{-}\Big{)}\to\Big{(}D^{+}\cup_{\mathrm{id}_{\partial D^{+}}}\big{(}\partial D^{+}\times[0,1]\big{)}\cup_{\varphi_{1}}D^{-}\Big{)}$
@@ 3. Proof of the main result
We may assume without loss of generality that $L^{+}\coloneq P^{-1}(\ell)$ is singular, i.e. of codimension $>1$. We will rely on the fact that any geodesic $\gamma$ in $M$ starting at $L^{0}$ with $\gamma^{\prime}(0)\perp L^{0}$ and $|\gamma^{\prime}(0)|=1$ is projected by $P$ to a piecewise geodesic in $[-l,l]$ parametrized by arclength whose break points occur only at $-l$ and $l$ (see [KL22, Propositions 3.2 and 12.5, Corollary 7.4]). Thus if $\gamma$ starts in the direction of $L^{+}$, then $\gamma((1+4k)\cdot\ell)\in L^{+}$ for all $k\geq 0$. Since $L^{+}$ has strictly smaller codimension than $L^{0}$, this implies that the points $\gamma((1+4k)\cdot\ell)$ for $k\geq 0$ are singular values for the normal exponential map $\exp\colon\nu_{g}(L^{0})\rightarrow M$, i.e. that they are focal points of $L^{0}$. The number of focal points for $L^{0}$ on $\gamma\colon(0,4n\ell)\rightarrow M$, counted with multiplicity, is thus at least $n$. Now suppose by contradiction that $\smash{4n\ell<{\tfrac{\pi}{\sqrt{\kappa}}}}$. It then follows from Rauch’s comparison theorem that $\gamma\colon(0,4n\ell)\rightarrow M$ has no conjugate points. But this implies by [Lyt09, Corollary 1.2] that the number of focal points of $L^{0}$ along $\gamma\colon(0,4n\ell)\rightarrow M$ is at most $\mathrm{dim}(L^{0})=n-1$ (this result is stated for manifolds without focal points, but the argument applies to any geodesic segment without conjugate points), so we arrive at a contradiction. ∎
@@
Equipped with this proposition, we now prove Theorem 3.1. Arguing by contradiction, we choose an infinite sequence $(M_{i},g_{i},P_{i})\in\mathcal{M}^{n,\kappa,r,\nu}$ of manifold submetries that are pairwise not foliated homeomorphic, and if $n\neq 5$ also pairwise not foliated diffeomorphic. By Lemma 2.3, the base $X_{i}$ is isometric to an interval, so up to postcomposition by an isometry, we may assume $X=[-l_{i},l_{i}]$ with $2l_{i}\leq\operatorname{diam}(M_{i})<r$. By Proposition 3.2, there is a common positive lower bound for all the $l_{i}$, so by rescaling by the uniformly bounded numbers $\frac{1}{l_{i}}$ and changing $\kappa,r,\nu$, we may assume $l_{i}=1$ so $X=[-1,1]$ for all $i$. By Cheeger–Gromov’s compactness theorem [Pet87, Theorem 4.4], upon choosing a subsequence we can find a smooth manifold $M$ and diffeomorphisms $\psi_{i}:M\to M_{i}$, such that the metrics $\psi_{i}^{*}(g_{i})$ $\mathcal{C}{}^{1}$-converge to a $\mathcal{C}{}^{1,1}$-metric $g_{\infty}$ on $M$. Moreover, the exponential maps $\exp_{i}\coloneqq(\psi_{i}^{-1}\circ\exp_{(M_{i},g_{i})}\circ d\psi_{i}):TM\to M$ converge uniformly on compact subsets to $\exp_{\infty}\coloneqq\exp_{(M,g_{\infty})}\colon TM\to M$. Identifying $M_{i}$ with $M$ via $\psi_{i}$, we may assume that $M_{i}=M$ and $\psi_{i}=\mathrm{id}_{M}$ for all $i$. By [KL22, Lemma 2.4], upon passing to a further subsequence, we can furthermore assume that the submetries $P_{i}:(M,g_{i})\rightarrow[-1,1]$ converge pointwise to a submetry $P_{\infty}:(M,g_{\infty})\to[-1,1]$ such that the fibers $P_{i}^{-1}(t)$ converge pointwise to the fibers $P_{\infty}^{-1}(t)$ for all $t$. The $\mathcal{C}{}^{1}$-Riemannian manifold $(M,g_{\infty})$ inherits the double-sided curvature bound in the sense of Alexandrov from the $(M,g_{i})$ (cf. [KL21]) and all fibers of $P_{\infty}$ are $\mathcal{C}^{1}$-submanifolds of $M$ by [KL22, Corollary 12.6]. We write
@@
Before turning to the proof of this proposition, we use it to finish the proof of Theorem 3.1. The derivatives of the $\alpha_{i}$ induce smooth bundle isomorphisms $d\alpha_{i}\colon\nu(\tilde{L}^{\pm})\rightarrow\nu(L_{i}^{\pm})$ between the normal bundles of the $\tilde{L}^{\pm}$ and $L_{i}^{\pm}$ as quotient vector bundles. Identifying the quotient bundle $\nu(L_{i}^{\pm})$ with the subbundle $\nu_{g_{i}}(L_{i}^{\pm})\subset TM$ using the Euclidean metric $g_{i}$ and pulling back $g_{i}$ along $d\alpha_{i}$, we obtain a sequence of smooth Euclidean metrics $d\alpha_{i}^{*}g_{i}$ on the vector bundle $\nu(\tilde{L}^{\pm})$ that converge against a $\mathcal{C}^{0}$-Euclidean metric $d\alpha_{\infty}^{*}g_{\infty}$. Now fix a smooth Euclidean metric $\tilde{g}$ on $\nu(\tilde{L}^{\pm})$ and choose smooth Euclidean bundle isomorphisms $u_{i}\colon(\nu(\tilde{L}^{\pm}),\tilde{g})\rightarrow(\nu(\tilde{L}^{\pm}),d\alpha_{i}^{*}g_{i})$ that cover the identity and converge uniformly on compact subsets against a $\mathcal{C}^{0}$ isomorphisms $u_{\infty}\colon(\nu(\tilde{L}^{\pm}),\tilde{g})\rightarrow(\nu(\tilde{L}^{\pm}),d\alpha_{\infty}^{*}g_{\infty})$ (this is possible since Euclidean bundles with respect to different metrics are isomorphic and these isomorphisms can be chosen to be close if the metrics are close). With respect to the resulting isomorphisms $w_{i}\coloneqq(u_{i}\circ d\alpha_{i})\colon(\nu(\tilde{L}^{\pm}),\tilde{g})\rightarrow(\nu_{g_{i}}(L_{i}^{\pm}),g_{i})$, the gluing diffeomorphism (3.1) becomes a diffeomorphism between the sphere bundles of $(\nu(\tilde{L}^{\pm}),\tilde{g})$,
@@
To construct the desired diffeomorphisms $\alpha_{i}$, we first construct for all $i$ large enough $\mathcal{C}^{1}$-diffeomorphisms $\eta_{i}^{+}:M\to M$ that converge in the $\mathcal{C}^{1}$-topology to the identity, satisfy $\eta_{i}^{+}(L_{i}^{+})=L_{\infty}^{+}$ and have support in a small neighborhood of $L_{\infty}^{+}$. To this end, we first change coordinates: we fix a $\mathcal{C}^{1}$-diffeomorphism $\zeta:M\to M$ supported in an arbitrarily small neighborhood of $L_{\infty}^{+}$ such that $\bar{L}^{+}:=\zeta(L_{\infty}^{+})$ is smooth (see e.g. [Hir94, Theorem 3.6] for the fact that such a $\zeta$ exists). Then the $\mathcal{C}^{1}$-submanifolds $\bar{L}_{i}^{+}:=\zeta(L^{+}_{i})$ converge in the Hausdorff-sense to $\bar{L}^{+}$, their tangent spaces at points in $\bar{L}_{i}^{+}$ converge to the corresponding tangent space at points in $\bar{L}^{+}$, and the volumes of $\bar{L}_{i}^{+}$ converge to the volume of $\bar{L}^{+}$. Now we smoothly identify a small tubular neighborhood of $\bar{L}^{+}$ with the smooth normal bundle. It follows that the foot-point projection of $\bar{L}_{i}^{+}$ to $\bar{L}^{+}$ is a $\mathcal{C}^{1}$ immersion, hence a covering map for all large $i$. Since this projection is almost $1$-Lipschitz, and the volumes converge, we deduce that this projection is injective. Therefore, $\bar{L}_{i}^{+}$ can be uniquely represented as a $\mathcal{C}^{1}$-section $h_{i}$ of the normal bundle of $\bar{L}^{+}$. The convergence statement for the tangent spaces, implies that this sections $h_{i}$ converge to the 00-section in the $\mathcal{C}^{1}$-topology. Using the sections $h_{i}$, we find $\mathcal{C}^{1}$-diffeomorphisms of the normal bundle that send $h_{i}$ to the 00-section, are supported in a small neighbourhood of the 00-section, and converge in the $\mathcal{C}^{1}$-topology. Extending these diffeomorphisms by the identity outside of the tubular neighbourhood of $\bar{L}^{+}$ and changing back coordinates using $\zeta^{-1}$, we obtain diffeomorphisms $\eta_{i}^{+}$ as claimed. Applying the same reasoning to $L_{\infty}^{-}$ and composing the resulting diffeomorphisms, we find diffeomorphisms $\eta_{i}:M\to M$ which converge in the $\mathcal{C}^{1}$-topology to the identity and satisfy $\eta_{i}(L_{i}^{\pm})=L_{\infty}^{\pm}$. Finally, we postcompose by a choice of a $\mathcal{C}^{1}$-diffeomorphism supported in a neighborhood of $L^{\pm}$ and $L^{0}$ in order to move $L^{\pm}$ to smooth submanifolds $\tilde{L}^{\pm}$ of $M$ (using e.g. [Hir94, Theorem 3.6]). Taking the inverse of the resulting $\mathcal{C}^{1}$-diffeomorphism yields diffeomorphisms $\alpha_{i}$ as in the claim.∎

### v2

@@ 1. Introduction
    The reason why dimension $n=5$ in Theorem 1.1 is special is that our proof relies on the fact that for a closed (n1)(n-1)-manifold $L$ the forgetful morphism $\pi_{0}(\operatorname{Diff}(L))\to\pi_{0}(\operatorname{Homeo}(L))$ has finite kernel for $n\neq 5$ (see Lemma 5.6). Here $\pi_{0}(\operatorname{Diff}(L))$ denotes the group of smooth isotopy classes of diffeomorphisms and $\pi_{0}(\operatorname{Homeo}(L))$ the group of topological isotopy classes of homeomorphisms. The analogous statement for $n=5$ is known to be false (see Remark 5.7).
@@
    A leaf is singular if it has codimension >1>1 and regular if it has codimension $1$.
@@ Acknowledgments
This work originated from discussions following a colloquium talk by MR as part of the Research Training Group DFG 281869850 (RTG 2229), funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). The authors would like to thank the RTG for providing a fruitful scientific environment. A.L. would also like to thank the Hausdorff Research Institute of Mathematics in Bonn where this work was finalised during a stay of A.L. as part of the Trimester Program “Metric Analyis”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC-2047/1–390685813.
@@ 2.1. Transnormal functions, manifold submetries, and double disc bundles
The second alternative point of view is that of double disc bundles. To introduce it, we say that Euclidean disc bundle $D\rightarrow L$ is the unit disc bundle of a vector bundle equipped with a fiberwise metric over a closed manifold $L$ which may be disconnected and have components of different dimension. A double disc bundle decomposition of $M$ consists of two Euclidean disc bundles $D^{\pm}\rightarrow L^{\pm}$, a diffeomorphism $\varphi\colon\partial D^{+}\rightarrow\partial D^{-}$ between the boundaries of their total spaces, and a diffeomorphism $\tau\colon D^{+}\cup_{\varphi}D^{-}\rightarrow M$ from the manifold obtained by gluing $D^{\pm}$ along the boundaries using $\varphi$. Given such decomposition, there is a canonical map $P_{D^{\pm}}\colon M\rightarrow[-1,1]$ which is given as the composition of $\tau^{-1}$ with the map $D^{+}\cup_{\varphi}D^{-}$ that sends $e\in D^{\pm}$ to $\pm(\lvert e\rvert-1)$, using the fiberwise norm of $D^{\pm}$. The fibers over $\pm 1$ are the 0-sections $L^{\pm}\subset M$ and the fibers over $t\in(-1,1)$ are the concentric sphere subbundles of $D^{\pm}$.
@@
We first show that (ii) and (iii) are equivalent. Given a manifold submetry $P\colon(M,g)\rightarrow[c,d]$ as in (ii), we can without changing the fibers rescale the metrics so that c=1c=-1 and $d=1$ (up to an opportune translation of the codomain). Now consider the smooth submanifolds $L^{\pm}\coloneqq P^{-1}(\pm 1)$ and the normal exponential map $\exp^{\pm}\colon\nu_{g}(L^{\pm})\rightarrow M$ from the normal bundle $\nu_{g}(L^{\pm})\subset TM$ of $L^{\pm}$ with respect to $g$. Using that $P$ is a submetry and that geodesics in [1,1][-1,1] with length <2<2 starting from $\pm 1$ minimize distance, one sees that the restrictions $\exp^{\pm}\colon D_{g}(L^{\pm})\rightarrow M$ to the unit disc bundles $D_{g}(L^{\pm})\subset\nu_{g}(L^{\pm})$ are diffeomorphisms onto $P^{-1}([-1,0])$ and $P^{-1}([0,1])$ respectively (c.f. [KL22, Section 3.4]). Under this identification, the map $P\colon M\rightarrow[-1,1]$ corresponds to $D_{g}(L^{\pm})\rightarrow[-1,1],e\mapsto\pm(\lvert e\rvert-1)$, so we obtain a double disc bundle composition $\smash{D_{g}(L^{+})\cup_{\varphi}D_{g}(L^{-})\cong M}$ with $\smash{P_{D_{g}(L^{\pm})}=P}$ and gluing diffeomorphism $\varphi=(\exp^{-})^{-1}\circ\exp^{+}$ involving the diffeomorphism $\smash{\exp^{\pm}\colon\partial D_{g}(L^{\pm})\rightarrow P^{-1}(0)}$ given by restricting the normal exponential map to the unit sphere bundle. This shows that (ii) implies (iii). To show the converse, one constructs a suitable metric $g$ starting from a double disc bundle decomposition (see [GWY19, Lemma 3.1, Remark 3.2]).
@@
To prove that (i) and (ii) are equivalent, we use that the fibers of a transnormal function $f\colon M\rightarrow\mathbb{R}$ as in (i) are pairwise equidistant smooth submanifolds of $M$ by [Wan87]. This implies that there is a canonical metric on $f(M)$ that makes $f\colon M\rightarrow f(M)$ into a manifold submetry. By the mean value theorem, $f(M)$ is homeomorphic to a compact interval and thus also isometric to a compact interval with the standard metric. Postcomposing $f$ with this isometry yields a manifold submetry as in (ii) with the same fibers as $f$. Vice versa, given a manifold submetry $P:M\to[-c,d]$, we may assume [c,d]=[1,1][c,d]=[-1,1] by rescaling the metric. We choose a smooth and strictly monotonically inreasing function $h\colon[-1,1]\rightarrow\mathbb{R}$ with h(1)=1h(-1)=-1, $h(1)=1$, $h^{\prime}(t)>0$ on (1,1)(-1,1), and such that $h(t)=(t+1)^{2}-1$ near 1-1 and $h(t)=1-(t-1)^{2}$ near $1$. The function $f\coloneq(h\circ P)$ has the same fibers as $P$, so it suffices to check that it is transnormal. As observed above, we have $P(-)=\pm 1\mp d_{M}(-,L^{\pm})$. Moreover, we have $(P\circ\exp^{\pm})(v)=\pm 1\mp|v|$ for $|v|\leq 2$, so it follows that $f=h\circ P$ is a smooth function. To check that it satisfies the first equation in (1.1), note that the differential of $f$ vanishes exactly on $L^{\pm}$. Near $L^{\pm}$ the function $f$ satisfies $\|\nabla f(\exp^{\pm}(v))\|^{2}=\|v\|^{2}$, which is a linear function of $f$. Outside of $L^{\pm}$, we have $\|\nabla f\|^{2}=\|h^{\prime}(P(x))\|^{2}=\|h^{\prime}\circ h^{-1}\circ f\|^{2}$. From this one concludes that $f$ satisfies indeed the first equation in (1.1), so it is transnormal.∎
@@ 2.2. Connectedness of fibers and existence of singular fibers
    $\mathcal{F}$ arises as an isoparametric foliation on MMwith respect to some Riemannian metric on $g$ in the sense of the introduction.
@@ 2.3. Foliated diffeomorphism, concordance, and isotopy
It is convenient to replace $D^{+}\cup_{\varphi_{i}}D^{-}$ up to foliated diffeomorphism with $\smash{D^{+}\cup_{\mathrm{id}_{\partial D^{+}}}\big(\partial D^{+}\times[0,1]\big)\cup_{\varphi_{i}}D^{-}}$. Here the latter is equipped with the partition into submanifolds given as the fibers of the map to [1,2][-1,2] that sends $e\in D^{+}$ to $\lvert e\rvert-1$, $e\in D^{-}$ to $2-\lvert e\rvert$, and agrees with the projection to $[0,1]$ on the middle part. If the $\varphi_{i}$ are smoothly isotopic, then $\mathrm{id}_{\partial D^{+}}$ and $\varphi_{0}^{-1}\circ\varphi_{1}$ are smoothly isotopic. A choice of isotopy $\phi\colon\partial D^{+}\times[0,1]\to\partial D^{+}\times[0,1]$ between them induces a diffeomorphism
$\hskip-9.95863pt(\mathrm{id}_{D^{+}}\cup\phi\cup\mathrm{id}_{D^{-}})\colon\Big(D^{+}\cup_{\mathrm{id}_{\partial D^{+}}}\big(\partial D^{+}\times[0,1]\big)\cup_{\varphi_{0}}D^{-}\Big)\to\Big(D^{+}\cup_{\mathrm{id}_{\partial D^{+}}}\big(\partial D^{+}\times[0,1]\big)\cup_{\varphi_{1}}D^{-}\Big)$
@@ 3. Proof of the main result
We may assume without loss of generality that $L^{+}\coloneq P^{-1}(\ell)$ is singular, i.e. of codimension >1>1. We will rely on the fact that any geodesic $\gamma$ in $M$ starting at $L^{0}$ with $\gamma^{\prime}(0)\perp L^{0}$ and $|\gamma^{\prime}(0)|=1$ is projected by $P$ to a piecewise geodesic in [l,l][-l,l] parametrized by arclength whose break points occur only at l-l and $l$ (see [KL22, Propositions 3.2 and 12.5, Corollary 7.4]). Thus if $\gamma$ starts in the direction of $L^{+}$, then $\gamma((1+4k)\cdot\ell)\in L^{+}$ for all $k\geq 0$. Since $L^{+}$ has strictly smaller codimension than $L^{0}$, this implies that the points $\gamma((1+4k)\cdot\ell)$ for $k\geq 0$ are singular values for the normal exponential map $\exp\colon\nu_{g}(L^{0})\rightarrow M$, i.e. that they are focal points of $L^{0}$. The number of focal points for $L^{0}$ on $\gamma\colon(0,4n\ell)\rightarrow M$, counted with multiplicity, is thus at least $n$. Now suppose by contradiction that $\smash{4n\ell<{\tfrac{\pi}{\sqrt{\kappa}}}}$. It then follows from Rauch’s comparison theorem that $\gamma\colon(0,4n\ell)\rightarrow M$ has no conjugate points. But this implies by [Lyt09, Corollary 1.2] that the number of focal points of $L^{0}$ along $\gamma\colon(0,4n\ell)\rightarrow M$ is at most $\mathrm{dim}(L^{0})=n-1$ (this result is stated for manifolds without focal points, but the argument applies to any geodesic segment without conjugate points), so we arrive at a contradiction. ∎
@@
Equipped with this proposition, we now prove Theorem 3.1. Arguing by contradiction, we choose an infinite sequence $(M_{i},g_{i},P_{i})\in\mathcal{M}^{n,\kappa,r,\nu}$ of manifold submetries that are pairwise not foliated homeomorphic, and if $n\neq 5$ also pairwise not foliated diffeomorphic. By Lemma 2.3, the base $X_{i}$ is isometric to an interval, so up to postcomposition by an isometry, we may assume $X=[-l_{i},l_{i}]$ with $2l_{i}\leq\operatorname{diam}(M_{i})<r$. By Proposition 3.2, there is a common positive lower bound for all the $l_{i}$, so by rescaling by the uniformly bounded numbers $\frac{1}{l_{i}}$ and changing $\kappa,r,\nu$, we may assume $l_{i}=1$ so X=[1,1]X=[-1,1] for all $i$. By Cheeger–Gromov’s compactness theorem [Pet87, Theorem 4.4], upon choosing a subsequence we can find a smooth manifold $M$ and diffeomorphisms $\psi_{i}:M\to M_{i}$, such that the metrics $\psi_{i}^{*}(g_{i})$ $\mathcal{C}{}^{1}$-converge to a $\mathcal{C}{}^{1,1}$-metric $g_{\infty}$ on $M$. Moreover, the exponential maps $\exp_{i}\coloneqq(\psi_{i}^{-1}\circ\exp_{(M_{i},g_{i})}\circ d\psi_{i}):TM\to M$ converge uniformly on compact subsets to $\exp_{\infty}\coloneqq\exp_{(M,g_{\infty})}\colon TM\to M$. Identifying $M_{i}$ with $M$ via $\psi_{i}$, we may assume that $M_{i}=M$ and $\psi_{i}=\mathrm{id}_{M}$ for all $i$. By [KL22, Lemma 2.4], upon passing to a further subsequence, we can furthermore assume that the submetries $P_{i}:(M,g_{i})\rightarrow[-1,1]$ converge pointwise to a submetry $P_{\infty}:(M,g_{\infty})\to[-1,1]$ such that the fibers $P_{i}^{-1}(t)$ converge pointwise to the fibers $P_{\infty}^{-1}(t)$ for all $t$. The $\mathcal{C}{}^{1}$-Riemannian manifold $(M,g_{\infty})$ inherits the double-sided curvature bound in the sense of Alexandrov from the $(M,g_{i})$ (cf. [KL21]) and all fibers of $P_{\infty}$ are $\mathcal{C}^{1}$-submanifolds of $M$ by [KL22, Corollary 12.6]. We write
@@
[…]
@@
[…]